OpenStudy (anonymous):

if f(2)=5 and f(-1)=-3, find a and b in the function f(x)=ax+b

OpenStudy (dan815):

they gave u 2 points

OpenStudy (dan815):

(2,5) and (-1,-3)

OpenStudy (dan815):

can you find the slope given these 2 points?

OpenStudy (dan815):

slope is the change in y per every change in x

OpenStudy (dan815):

for the points (2,5) and (-1,-3) look at how the y changes and how the x changed y from 5 to -3 x from 2 to -1

OpenStudy (anonymous):

the slope would be -2 /-3

OpenStudy (dan815):

y went down -8 x went down -3 -8/-3 = 8/3

OpenStudy (dan815):

remember from 5 to -3 not 5 to 3

OpenStudy (dan815):

plug in your slope and one of those points and see what b has to be

OpenStudy (anonymous):

oh yes i got -8/-3

OpenStudy (dan815):

f(x)=ax+b f(x)=8/3x+b pick one of the points (2,5) 5=8/3*(2)+b b=?

OpenStudy (dan815):

f(x) is another way to say y

OpenStudy (dan815):

y=8/3x+b we are given point (2,5)<---- x,y value in that order so we can plug that in for x,y 5=8/3(2)+b so we can see what b has to be so that this equation is satisfied

OpenStudy (anonymous):

i got -1/3

OpenStudy (dan815):

yes

OpenStudy (anonymous):

thank you so much, i appreciate it

OpenStudy (dan815):

okay sure thing :)